A bicycle dealer was asked how many bikes he had in stock. He answered, “If one-half,

one-third, and one-quarter of the number of bikes were added together, they would make

13.” How many bikes did he have in stock?



1
Expert's answer
2021-10-18T15:19:23-0400

Solution:

Let the total number of bikes in stock be "x".

Now, according to question,

"\dfrac 12x+\dfrac 13x+\dfrac 14x=13"

"\Rightarrow x(\dfrac {6+4+3}{12})=13
\\\Rightarrow x(\dfrac {13}{12})=13
\\\Rightarrow x=13\times (\dfrac {12}{13})
\\\Rightarrow x=12"